OpenStudy (anonymous):

Four points are placed on a circle. How many segments will it take to connect each point to every other point?

OpenStudy (anonymous):



OpenStudy (anonymous):

This is same as constructing a complete graph with 4 vertices. \[\frac{(4 \times 3)}{2} = 6 \]

OpenStudy (anonymous):

Or doe the circle itself resembles a connection?

OpenStudy (anonymous):

If that is the case then the answer is 2.

OpenStudy (anonymous):

Assume you are referring to line segments. Little bit of experimenting will show you that each time you connect a dot to the remaining it reduces by 1 so 3+2+1 = 6